Pappy's Toys makes two models of a metal toy—Standard and DeLuxe. Both models are produced on a single machine. The price and costs of the two models appear as follows.

Standard

DeLuxe

Price per unit

$

140

$

160

Variable cost per unit

$

60

$

64

Machine hours per unit

0.5

1.5

The one machine that used to produce both models has a capacity of 110,000 machine-hours per quarter. Fixed manufacturing costs per quarter are $900,000.

Required:

a. Suppose that the maximum unit sales in a quarter that Pappy's can achieve is 240,000 units of each product. How many units of each model should Pappy's produce in a quarter?

b. Suppose that the maximum unit sales in a quarter that Pappy's can achieve is 130,000 units of each product. How many units of each model should Pappy's produce in a quarter?

A.

Standard Model

Units

Deluxe Model

Units

B.

Standard Model

Units

Deluxe Model

Units
Answer #1

A) standard model 110,000/0.5= 220,000

deluxe model: 0


B) standard model: 130,000

deluxe model: 43,333

130,000*0.5= 65,000

130,000-65,000=65,000

65,000/1.5= 43,333.333333333 

answered by: a
Answer #2
Working Notes: 1
Calculation of Contribution margin per unit of Machine hours and ranking of the same
Product - Standard Product - Deluxe
Price Per Unit $                              140 $                           160
Less: Variable Cost Per Unit $                                 60 $                             64
Contribution Margin $                                 80 $                             96
Divide By "/" By "/" By
Machine hours required per unit of production 0.5 1.5
Equals to = (1 Hrs / 4 Units) (2 Hrs / 4 Units)
Contribution Margin Per machine hours $                           160.0 $                          64.0
Ranking of the same 1st   2 Nd
Solution: A
Calculation of optimum production for maximum contribution margin if 240,000 units sales available
Units Machine hrs required Total hrs used Balane Hrs Avaialbe
Total Hrs Avaialble 110000
Product - Standard 220000 0.5 110000 0
If the Maximum unit sales is 240,000 for the both units than,
Pappy casn produce only Standard 220,000 units
Answer =
Standard Model 220000 Units
Deluxe Model 0 Units
Solution: B
Calculation of optimum production for maximum contribution margin if 130,000 units sales available
Units Machine hrs required Total hrs used Balane Hrs Avaialbe
Total Hrs Avaialble 110000
Product - Standard 130000 0.5 65000 45000
Product - Delux   30000 1.5 45000
(45,000 Hrs / 1.5 Hrs)
Answer =
Standard Model 130000 Units
Deluxe Model 30000 Units